# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

### v3

@@ III Benchmarking LLM4AD in Simulation
This section introduces two comprehensive benchmarks for evaluating the instruction-following capabilities and performance of LLM-based agents in autonomous driving: the LaMPilot-Bench dataset on the HighwayEnv simulator and the CARLA Leaderboard 1.0 [dosovitskiy_carla_2017].
III-A LaMPilot-Bench
III-A1 Dataset
@@
The LaMPilot-Bench dataset consists of 4.9K semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes a high-level task description (instruction), an initial state for initializing the simulator, and goal state criteria aligned with the instruction. The dataset covers diverse driving scenarios, as shown in Tab. II. For each scenario, LaMPilot-Bench includes various situations, such as different ego vehicle initial positions and states, specific tasks (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed.
TABLE II: Main statistics of the LaMPilot-Bench dataset.
@@
The LaMPilot-Bench simulator builds upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making, which offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
III-A2 Evaluator
The LaMPilot-Bench evaluator incorporates metrics to assess the safety and efficiency of the agent driving policies. Time-to-collision (TTC) measures the vehicle’s ability to maintain safe distances and avoid collisions. In a scenario with $n+1$ vehicles (ego vehicle labeled as 0), the TTC ($\tau_{i}$) with vehicle $i$ is calculated using the ego vehicle’s longitudinal velocity $\mathbf{v}_{0}$ and position $\mathbf{p}_{0}$, and vehicle $i$’s longitudinal velocity $\mathbf{v}_{i}$ and position $\mathbf{p}_{i}$ ($1\leq i\leq n$):
$\tau_{\min}=\min_{1\leq i\leq n,1\leq t\leq T}\tau^{t}_{i}$ (2)
where $\mu_{0}$ is the mean speed:
$\mu_{0}=\frac{\sum_{t=1}^{T}{\|\mathbf{v}^{t}_{0}\|}}{T}.$ (5)
The SV score is defined as:
III-B CARLA Leaderboard 1.0 Benchmark
In addition to LaMPilot-Bench and HighwayEnv, we also utilize the CARLA Leaderboard 1.0 benchmark [carla_team_autonomous_2020, dosovitskiy_carla_2017] to evaluate LLM-based agents in a more realistic and complex driving environment. The Leaderboard consists of 76 routes (50 training, 26 testing) totaling over 170 km across six towns. It also includes predefined scenarios covering ten types of challenging traffic situations, such as control loss and traffic negotiation. To test the instruction-following capabilities of LLM-powered systems, we modify the Leaderboard setup to provide natural language navigation instructions based on the agent’s current position instead of GPS coordinates. To eliminate perception influence on our results, we craft the driving context information using a language generator that leverages CARLA’s privileged information. Executing an LMP in CARLA involves using Python’s exec function, which takes the LMP code as an input string and two dictionaries defining the execution scope: (i) apis, containing all driving APIs the code may call, and (ii) local_vars, initially empty but later holding a generator variable named policy once executed. The Leaderboard evaluates agent performance using three key metrics:
III-C Baselines and Results
We establish several baselines to evaluate the performance of LLM-based agents across the benchmarks, as presented in Tab. III for LaMPilot-Bench and Tab. IV for the CARLA Leaderboard 1.0.
TABLE III: Comparison of baselines on LaMPilot-Bench under different learning settings.
Method Learning Collision ($\downarrow$) Completion ()($\uparrow$)( ↑ ) TTC Score ()($\uparrow$)( ↑ ) SV Score ()($\uparrow$)( ↑ ) TE Score ()($\uparrow$)( ↑ ) Driving Score ()($\uparrow$)( ↑ )
@@ III-C Baselines and Results
TABLE IV: Driving performance of expert drivers on CARLA leaderboard 1.0 testing routes. $\uparrow$: Higher values are better.
Method Learning Navigation DS ()($\uparrow$)( ↑ ) RC ()($\uparrow$)( ↑ ) IP ()($\uparrow$)( ↑ )
@@
Rule-based methods such as the Intelligent Driver Model (IDM)[treiber_congested_2000] and Minimizing Overall Braking Induced by Lane Changes (MOBIL)[kesting_general_2007]. Both methods achieve a zero collision rate in LaMPilot-Bench, as shown in Tab. III. However, it is important to note that these methods operate independently of the provided instructions. Without considering human instructions, IDM and MOBIL achieve success rates of 20.4% and 15.3%, respectively. These results provide a reference point for assessing the effectiveness of LLM-based agents in following human instructions.
@@
LLMs are prompted with task instructions and API documentation, with or without in-context examples. For the few-shot setting, human programmers proficient in Python create in-context examples. They are given API documentation and allowed to write and test their code on the training routes. The same examples are used across all test routes. In the zero-shot setting, GPT models [openai_gpt-4_2023] and PaLM 2 [anil_palm_2023], given only the API Docs, driving context, and instructions, achieve an obvious performance advantage compared to rule-based methods in terms of task completion, as shown in Tab. III. When provided with training examples that include exemplar code (three-shot setting), all the evaluated LLMs exhibit notable improvements in completion rates. For instance, Llama 2 [touvron_llama_2023]’s completion rate increases from 20.4% in the zero-shot setting to 63.3% in the three-shot setting.
@@
LLMs are integrated with a human-in-the-loop learning approach, leveraging human feedback to refine generated policies. After executing a generated policy code ($P$), the human evaluators provide natural language feedback ($F$), which is fed back into the LLM along with $P$. This feedback loop enables continuous learning. If the feedback is positive (i.e., the human is satisfied with the execution), the code ($P$) is committed to the database for future retrieval and reuse. Otherwise, the feedback serves as guidance for iterative improvement. The integration of GPT models with human-in-the-loop learning further enhances their performance in driving tasks. Notably, GPT-4 achieves the highest driving score of $64.0$, as shown in Tab. III. This improvement highlights the great potential of LLMs in following instructions in the driving context when combined with human feedback.
@@
Comparison with other methods
We compare the human-feedback baseline with reinforcement learning experts that have access to privileged information, such as Roach [zhang_end--end_2021] and TCP [wu_trajectory-guided_2022] in the CARLA leaderboard [ma_learning_2024], as shown in Tab. IV. The comparison aims to provide a reference to previous works. Based on these results, we observe the following:
@@
The experiments demonstrate that off-the-shelf LLMs can generate code policies for driving tasks based on human instructions, outperforming rule-based methods in task completion. Few-shot learning with exemplar code further improves the performance of LLM-based agents, and the integration of human feedback enables iterative refinement of generated policies, leading to the highest driving scores among the evaluated methods. However, the notable collision rates in both benchmarks indicate the need for further research to fully capture the complexities and safety requirements of real-world driving scenarios. Most failures stem from the trade-off made for the LLM low-level planning frequency, as illustrated in the example below:

### v4

@@ IV Benchmarking LLM4AD in Simulation
This section introduces two comprehensive benchmarks for evaluating the instruction-following capabilities and performance of LLM-based agents in autonomous driving: One is for LLM-based agent planning evaluation (the LaMPilot-Bench dataset on the HighwayEnv simulator and the CARLA Leaderboard 1.0 [dosovitskiy_carla_2017]); Another is a multi-view visual question answering benchmark (NuPlanQA based on Nuplan dataset [caesar2022nuplanclosedloopmlbasedplanning, park2025nuplanqalargescaledatasetbenchmark]).
IV-A LaMPilot-Bench in HighwayEnv
IV-A1 Dataset
@@
The LaMPilot-Bench dataset consists of 4.9K semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes a high-level task description (instruction), an initial state for initializing the simulator, and goal state criteria aligned with the instruction. The dataset covers diverse driving scenarios, as shown in Tab. III. Our “semi-human-annotated” protocol involved a two-stage process. First, for each scenario, LaMPilot-Bench includes various situations, such as different ego vehicle initial positions and states, specific tasks (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed. Second, these scenarios were presented to a team of human annotators. These annotators were tasked with writing a clear, natural language instruction (e.g., “Move to the middle lane and drive at 30 mph”) that corresponded to the programmatic goal. To ensure the quality and consistency of these annotations, we then implemented a quality control pass. A separate team of senior evaluators reviewed the annotations, checking for ambiguity, grammatical errors, and, most importantly, a clear alignment between the natural language instruction and the scenario’s programmatic goal. Any instruction that was found to be ambiguous or did not uniquely map to the intended goal was discarded and sent back for re-annotation.
The LaMPilot-Bench simulator is built upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making, which offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
IV-A2 Evaluator
The LaMPilot-Bench evaluator incorporates metrics to assess the safety and efficiency of the agent driving policies. Time-to-collision (TTC) measures the vehicle’s ability to maintain safe distances and avoid collisions. In a scenario with $n+1$ vehicles (ego vehicle labeled as 0), the TTC with vehicle $i$, denoted as $\tau_{i}$, is calculated using the ego vehicle’s longitudinal velocity $\mathbf{v}_{0}$ and position $\mathbf{p}_{0}$, and vehicle $i$’s longitudinal velocity $\mathbf{v}_{i}$ and position $\mathbf{p}_{i}$ ($1\leq i\leq n$):
$\tau_{\min}=\min_{1\leq i\leq n,1\leq t\leq T}\tau^{t}_{i}$ (2)
The expression $100-1/\tau_{min}$ is a hyperbolic penalty function. This form is chosen because the penalty $1/\tau_{min}$ grows non-linearly toward infinity as $\tau_{min}$ approaches the unsafe boundary of 0. This ensures that critically dangerous, near-collision events are penalized far more severely than marginally unsafe ones. This “inverse barrier” logic is a common method for enforcing safety constraints in robotics and is a core concept in Control Barrier Functions (CBFs) [ames2016control].
where $\mu_{0}$ is the mean speed:
$\mu_{0}=\frac{\sum_{t=1}^{T}{\|\mathbf{v}^{t}_{0}\|}}{T}.$ (5)
The SV score is defined as:
IV-A3 Baselines and Results
We establish several baselines to evaluate the performance of LLM-based agents across the benchmarks, as presented in Tab. IV for LaMPilot-Bench.
TABLE III: Main statistics of the LaMPilot-Bench dataset.
@@ IV-A3 Baselines and Results
TABLE IV: Comparison of baselines on LaMPilot-Bench under different learning settings.
Method Learning Collision ($\downarrow$) Completion $(\uparrow)$ TTC Score $(\uparrow)$ SV Score $(\uparrow)$ TE Score $(\uparrow)$ Driving Score $(\uparrow)$
@@
TABLE V: Driving performance of expert drivers on CARLA leaderboard 1.0 testing routes. $\uparrow$: Higher values are better.
Method Learning Navigation DS $(\uparrow)$ RC $(\uparrow)$ IP $(\uparrow)$
@@
Rule-based methods such as the Intelligent Driver Model (IDM)[treiber_congested_2000] and Minimizing Overall Braking Induced by Lane Changes (MOBIL)[kesting_general_2007]. Both methods achieve a zero-collision rate in LaMPilot-Bench, as shown in Tab. IV. However, it is important to note that these methods operate independently of the provided instructions. Without considering human instructions, IDM and MOBIL achieve success rates of 20.4% and 15.3%, respectively. These results provide a reference point for assessing the effectiveness of LLM-based agents in following human instructions.
@@
LLMs are prompted with task instructions and API documentation, with or without in-context examples. For the few-shot setting, human programmers proficient in Python create in-context examples. They are given API documentation and allowed to write and test their code on the training routes. The same examples are used across all test routes. In the zero-shot setting, GPT models [openai_gpt-4_2023] and PaLM 2 [anil_palm_2023], given only the API Docs, driving context, and instructions, achieve an obvious performance advantage compared to rule-based methods in terms of task completion, as shown in Tab. IV. When provided with training examples that include exemplar code (three-shot setting), all the evaluated LLMs exhibit notable improvements in completion rates. For instance, Llama 2 [touvron_llama_2023]’s completion rate increases from 20.4% in the zero-shot setting to 63.3% in the three-shot setting.
@@
LLMs are integrated with a human-in-the-loop learning approach, leveraging human feedback to refine generated policies. After executing a generated policy code ($P$), the human evaluators provide natural language feedback ($F$), which is fed back into the LLM along with $P$. This feedback loop enables continuous learning. If the feedback is positive (i.e., the human is satisfied with the execution), the code ($P$) is committed to the database for future retrieval and reuse. Otherwise, the feedback serves as guidance for iterative improvement. The integration of GPT models with human-in-the-loop learning further enhances their performance in driving tasks. Notably, GPT-4 achieves the highest driving score of $64.0$, as shown in Tab. IV. This improvement highlights the great potential of LLMs in following instructions in the driving context when combined with human feedback.
@@
IV-B CARLA Leaderboard 1.0 Benchmark
IV-B1 Dataset
In addition to LaMPilot-Bench in HighwayEnv, we also utilize the CARLA Leaderboard 1.0 benchmark [carla_team_autonomous_2020, dosovitskiy_carla_2017] to evaluate LLM-based agents in a more realistic and complex driving environment. The Leaderboard consists of 76 routes (50 for training, 26 for testing), totaling over 170 km across six towns. It also includes predefined scenarios covering ten types of challenging traffic situations, such as control loss and traffic negotiation. To test the instruction-following capabilities of LLM-powered systems, we modify the Leaderboard setup to provide natural language navigation instructions based on the agent’s current position instead of GPS coordinates. To eliminate perception influence on our results, we craft the driving context information using a language generator that leverages CARLA’s privileged information. Executing an LMP in CARLA involves using Python’s exec function, which takes the LMP code as an input string and two dictionaries defining the execution scope: (i) apis, containing all driving APIs the code may call, and (ii) local_vars, initially empty but later holding a generator variable named policy once executed. The Leaderboard evaluates agent performance using three key metrics:
IV-B2 Baselines and Results
Tab. V is for the CARLA Leaderboard 1.0. It is important to note that the scores reported in Tab. V represent the mean performance of each agent aggregated across all 26 official CARLA Leaderboard 1.0 testing routes with different random seeds.
Figure 4: An overview of NuPlanQA. NuPlanQA comprises nine subtasks across three skill areas to support the context-aware analysis of traffic scenes. The proposed baseline, BEV-LLM, is trained on NuPlanQA-1M using historical frames, BEV features from multi-view images, and control signals as inputs. Finally, we evaluate MLLMs using NuPlanQA-Eval, a multiple-choice QA benchmark for driving scene understanding.
We compare the human-feedback baseline with reinforcement learning experts that have access to privileged information, such as Roach [zhang_end--end_2021] and TCP [wu_trajectory-guided_2022] in the CARLA leaderboard [ma_learning_2024], as shown in Tab. V. The comparison aims to provide a reference to previous works. Based on these results, we observe the following:
@@ IV-B2 Baselines and Results
The experiments demonstrate that off-the-shelf LLMs can generate code policies for driving tasks based on human instructions, outperforming rule-based methods in task completion. Few-shot learning with exemplar code further improves the performance of LLM-based agents, and the integration of human feedback enables iterative refinement of generated policies, leading to the highest driving scores among the evaluated methods.
IV-C NuPlanQA Benchmark
IV-C1 Dataset
While LaMPilot-Bench and CARLA Leaderboard evaluate LLM-based agents in closed-loop simulation, they are limited to low-level control policies and rule-based task execution. To further assess the high-level reasoning and perception capabilities of MLLMs in realistic driving environments, we incorporate the recently proposed NuPlanQA benchmark [park2025nuplanqalargescaledatasetbenchmark].
NuPlanQA provides a large-scale, multi-view VQA framework tailored for driving-scene understanding. It is built upon the nuPlan [caesar2022nuplanclosedloopmlbasedplanning] dataset, offering 1M VQA pairs (NuPlanQA-1M) for training and 8K multiple-choice VQA pairs (NuPlanQA-Eval) for evaluation. Each VQA sample corresponds to multi-view frames with annotated control signals and scene descriptions. The dataset supports context-aware reasoning over both spatial and temporal dimensions of traffic scenes, bridging the gap between low-level control evaluation (as in LaMPilot-Bench) and high-level situational reasoning.
TABLE VI: Performance of MLLMs on NuPlanQA-Eval. The metric used is accuracy (%)—calculated as the number of correct responses over the total number of questions. The best-performing model in each task is bolded, while the second-best is underlined. LLaVA-OV: LLaVA-OneVision, LLaVA-NV: LLaVA-Next-Video.
Road Env. Perception Spatial Relations Recog. Ego-centric Reasoning
Method Trfc. Light Wea -ther Road Type Avg. Sur. Obj. Trfc. Flow Key Obj. Avg. Ego Ctrl. Situ. Asse. Act. Rec. Avg. Total
Multi-frame as Input
GPT-4o [openai2024gpt4ocard] 68.5 91.7 95.0 85.1 79.6 76.5 67.7 74.6 86.8 85.1 82.4 84.8 81.5
Gemini-1.5-Pro [gemini_team_gemini_2024] 64.5 93.5 95.8 84.6 69.1 78.6 73.2 73.6 73.8 73.8 80.4 76.0 78.1
VideoLLaMA27B[damonlpsg2024videollama2] 54.2 59.4 93.8 69.1 70.7 78.6 66.4 71.9 72.3 82.7 81.4 78.8 73.3
Qwen2.5-VL7B [bai2025qwen25vltechnicalreport] 51.7 71.9 69.3 64.3 49.2 23.0 30.9 34.4 30.4 39.6 10.1 26.7 41.8
LLaVA-OV7B [li2024llavaonevisioneasyvisualtask] 53.2 89.4 96.4 79.7 77.9 77.6 73.2 74.3 75.4 76.2 82.9 77.5 77.8
LLaVA-NV7B [zhang2024llavanext-video] 44.3 46.5 72.4 54.4 57.5 52.0 44.1 51.2 57.6 65.8 71.4 64.9 56.8
InternVL-1.520B [internvl1.5] 40.9 70.5 75.5 62.3 48.6 50.0 44.5 47.7 57.1 55.0 60.8 57.6 55.9
LLaVA-NV32B [zhang2024llavanext-video] 47.8 74.7 93.8 72.1 69.1 74.0 63.6 68.9 72.8 69.3 76.9 73.0 71.3
BEV-LLM (Ours) 61.1 89.4 89.6 80.0 78.5 75.5 68.2 74.1 79.1 83.2 83.4 81.9 78.7
Single-frame as Input
GPT-4o [openai2024gpt4ocard] 66.5 92.2 94.4 84.4 81.2 77.0 68.5 75.6 86.4 86.6 81.4 84.8 81.6
Gemini-1.5-Pro [gemini_team_gemini_2024] 64.5 93.1 95.8 84.5 71.1 79.6 72.7 74.5 79.6 76.7 80.9 79.1 79.4
VideoLLaMA27B[damonlpsg2024videollama2] 50.2 64.1 93.8 69.4 71.3 79.6 64.1 71.7 69.6 84.2 81.4 78.4 73.2
Qwen2.5-VL7B [bai2025qwen25vltechnicalreport] 51.2 70.5 62.5 61.4 47.0 16.8 23.2 29.0 27.2 40.6 12.1 26.6 39.0
LLaVA-OV7B [li2024llavaonevisioneasyvisualtask] 52.2 93.5 96.4 80.7 72.9 74.0 68.6 71.8 72.3 72.3 79.9 74.8 75.8
LLaVA-NV7B [zhang2024llavanext-video] 48.8 41.5 63.5 51.3 51.9 50.5 41.8 48.1 56.5 60.4 69.8 62.2 53.9
InternVL-1.520B [internvl1.5] 39.4 68.7 72.4 60.2 54.7 48.5 41.4 48.2 62.3 58.9 58.8 60.0 56.1
LLaVA-NV32B [zhang2024llavanext-video] 46.8 73.3 91.7 70.6 70.7 71.9 58.2 66.9 71.2 68.8 82.9 74.3 70.6
BEV-LLM (Ours) 58.1 87.1 84.9 76.7 75.1 75.0 66.8 72.3 76.0 81.2 81.4 79.5 76.2
Figure 5: Architecture of BEV-LLM. BEV features are integrated into LLaMA via the BEV-Fusion module.
As shown in Fig. 4, NuPlanQA decomposes the driving-scene understanding problem into nine subtasks spanning three hierarchical skill domains:
  • Road Environment Perception: traffic light recognition, weather/lighting condition, and road type/condition;
  • Spatial Relations Recognition: identification of surrounding objects, traffic flow, and key objects critical to the ego vehicle;
  • Ego-Centric Reasoning: inference of ego-vehicle maneuver, situation assessment, and action recommendation.
This categorization enables fine-grained evaluation of MLLMs across low-level perception, spatial understanding, and high-level reasoning, facilitating diagnostic analysis of LLMs’ strengths and weaknesses.
Each question in NuPlanQA-Eval is associated with four answer choices and evaluated by accuracy. To ensure annotation quality, all QA pairs undergo both GPT-4o-assisted generation with chain-of-thought prompting and human verification for semantic consistency and contextual correctness. The evaluation set maintains an approximately uniform distribution across all nine subtasks to prevent bias toward particular skill domains.
IV-C2 NuplanQA Baseline and Results
To establish a reference model, BEV-LLM is proposed as a multi-view, Bird’s-Eye-View(BEV)-enhanced MLLM that integrates spatially aligned BEV features from multiple cameras into the vision encoder of LLaMA-3.2-Vision [grattafiori2024llama3herdmodels]. As shown in Fig. 5, the LLaMA-3.2-Vision encoder first extracts per-frame visual features $\hat{V}_{t}$ from time-sequential multi-view images. The features across all timesteps are denoted as $V_{t}$.
To capture temporal dependencies, BEV features ($B_{t}$) are constructed by attending to historical BEV embeddings ($B_{t-1}$) using the temporal self-attention (TSA) from BEVFormer [li_bevformer_2022]:
$B_{\text{t}}=\text{TSA}(Q,\{Q,B_{t-1}\})+B_{t-1}$ (9)
where $Q$ represents BEV queries. This TSA mechanism encodes motion-aware scene representations by incorporating past information.
These refined BEV features are then fused with the multi-view visual features ($\hat{V}_{T}$) using spatio-temporal fusion attention (SFA). In SFA, visual features act as queries, while the temporally-enriched BEV embeddings ($B_{\text{proj}}$, projected from $B_{t}$) serve as keys and values:
$\hat{V}_{\text{BEV-att}}=\text{SFA}(\hat{V}_{T},B_{\text{proj}})+\hat{V}_{T}$ (10)
This cross-attention allows the visual features to integrate the spatial-temporal BEV context, implicitly transferring temporal awareness.
We extract features from multiple layers of $\hat{V}_{\text{BEV-att}}$ (every fourth layer and the final hidden state [grattafiori2024llama3herdmodels]) and concatenate them to create a robust multi-scale representation. A multi-modal projector then aligns this representation with the language model’s text embedding space.
Finally, the projected vision-BEV embeddings are fed into the LLaMA decoder alongside tokenized text instructions and vehicle data (e.g., velocity, steering). The decoder’s cross-attention layers function as a multi-modal adapter [grattafiori2024llama3herdmodels, alayrac2022flamingovisuallanguagemodel], enabling BEV-LLM to integrate motion, depth, and control signals for a comprehensive multi-modal understanding.
As shown in Tab. VI, on NuPlanQA-Eval, BEV-LLM achieves an average accuracy of 78.7%, outperforming other open-source models in six of nine subtasks, particularly in ego-centric reasoning and traffic light detection. These results highlight the effectiveness of BEV integration in enabling models to reason about motion and ego-vehicle context.
IV-D Summary
The LaMPilot-Bench dataset and the CARLA Leaderboard 1.0 benchmark provide a comprehensive evaluation framework for assessing the instruction-following capabilities and performance of LLM-based agents in autonomous driving. The experiments highlight the potential of LLMs in enhancing the decision-making capabilities of autonomous vehicles when combined with human feedback and domain-specific APIs. As LLMs continue to evolve and integrate with more modalities, we envision their significant contribution to the development of safe, efficient, and user-friendly autonomous vehicles. Future research should focus on addressing the challenges of real-time decision-making, reducing collision rates, and incorporating multi-modal inputs to further enhance the performance and robustness of LLM-based autonomous driving systems.
However, the notable collision rates in both benchmark indicate the need for further research to fully capture the complexities and safety requirements of real-world driving scenarios. Most failures stem from the trade-off made for the LLM low-level planning frequency, as illustrated in the example below: